# From MTEB to MTOB: Retrieval-Augmented Classification for Descriptive Grammars

### v1

@@ Abstract
Recent advances in language modeling have demonstrated significant improvements in zero-shot capabilities, including in-context learning, instruction following, and machine translation for extremely under-resourced languages (Tanzer et al., 2024). However, many languages with limited written resources rely primarily on formal descriptions of grammar and vocabulary.
@@
This set of benchmarks offers the first comprehensive evaluation of language models’ in-context ability to accurately interpret and extract linguistic features, providing a critical resource for scaling NLP to low-resource languages. The code and data are publicly available at https://anonymous.4open.science/r/from-MTEB-to-MTOB-6B0F.
@@ 1 Introduction
The advent of text-based foundational models has accelerated advancements in natural language processing, enhancing multilingual capabilities and applied tasks such as zero-shot machine translation, reading comprehension, and information extraction. Innovations like Machine Translation from One Book (MTOB) (Tanzer et al., 2024) utilize descriptive grammars to improve translation performance for extremely low-resource languages, showing the potential of large-scale language models in bridging formal linguistic theory and practical NLP applications. Despite the promise, using descriptive grammars for zero-shot MT presents challenges like terminology variability, non-standard structures, and scattered relevant information.
@@
  1. 1.
    The first scaled linguistic evaluation of the LLM machine reading capabilities on descriptive grammars.
  2. 2.
    A pipeline based on Retrieval Augmented Generation (RAG), which extracts relevant paragraphs from grammars based on a given typological characteristic (for example, WALS 81A: Order of Subject, Object and Verb [^1] https://wals.info/feature/81A) and provides them as prompts to an LLM to determine the meaning of these characteristics (for example, Subject-Verb-Object). The pipeline is evaluated through extensive experiments.
  3. 3.
    A benchmark consisting of 700 paragraphs from 14 descriptive grammars, annotated according to whether a linguist can unambiguously determine information about word order (WALS 81A) in the language described, in order to evaluate the quality of information retrieval methods on the task of filtering out the irrelevant paragraphs separately from the RAG pipeline.
  4. 4.
    A benchmark for the RAG pipeline, consisting of 148 grammars for each feature, in order to assess LLMs’ capabilities of determining typological characteristics based on the entire grammar at once and evaluate the effectiveness of different combinations of RAG pipeline components (i.e., information retrieval methods and prompts).
The proposed framework, alongside the presented benchmarks, aims to contribute to the ongoing efforts to improve the quality and efficiency of machine translation systems and to aid linguists in typological research by semi-automating extraction of data from descriptive grammars and. The entire codebase is open-source, licensed under MIT, and available at https://anonymous.4open.science/r/from-MTEB-to-MTOB-6B0F, promoting transparency and reproducibility within the community.
@@ 2 Related Work
Several advances in language modeling outline the possibility to leverage formal linguistic descriptions: retrieval-augmented generation methods for information extraction and generation, in-context abilities to operate with the formal texts in the prompt, and the availability of the formal texts in the machine-readable format.
@@ 2.1 Retrieval-Augmented Generation
RAG’s potential for handling low-resource languages and complex formal descriptions, such as linguistic grammars, has not been fully explored. Recent advances in retrieval-based frameworks point toward its applicability in linguistic resource-scarce domains, where available data is often fragmented and incomplete (Gao et al., 2024). This paper leverages the principles of RAG to enhance machine translation and language modeling for under-resourced languages through the use of descriptive grammars.
@@ 2.2 Zero-shot Learning
However, current zero-shot models still struggle with languages that have extremely limited or no monolingual or bilingual corpora. Recent studies, such as (Tanzer et al., 2024) and (Zhang et al., 2024), have shown that integrating formal linguistic descriptions—such as those found in grammars—can significantly improve zero-shot performance for these languages. In this work, we aim to further this line of research by evaluating the ability of models to utilize descriptive grammars in zero-shot settings.
@@ 4.1 Data
Examples of paragraphs for every relevance category are provided in Appendix A.
Among as few as 700 paragraphs, 38.86% (annotated with 0) are not relevant to order of any elements in a language. Furthermore, 42.43% of the paragraphs (the ones annotated with 1, 2, 3) are potentially misleading data due to describing order of components in a language without explicitly stating the value of WALS 81A. Finally, 18.71% paragraphs provide relevant information; however, out of the relevant paragraphs, 31.29% are only partially relevant (annotated with 4), and do not sufficiently elaborate on the constituent order in the monotransitive construction in order for the linguist to be able to determine the value of the feature. More detailed data on the benchmark is provided in Appendix B.
@@ 4.2 Results: Evaluating Rerankers
Wiki Summary 0.7474 0.7601 0.6535 0.7520 0.7690 0.7676 0.7776
Table 1: NDCG@20 scores on the reranker benchmark. The best result for each model is shown in bold, and the best result across all configuration variations is underlined. The NDCG@20 score for BM25 without a reranker is 0.7494.
The rerankers we chose are the 6 models with the best NDCG@10 score on the Massive Text Embedding Benchmark (MTEB) leaderboard (Muennighoff et al., 2023) on the Retrieval task for English as of Aug 29, 2024, along with SFR-Embedding-Mistral [^2] https://huggingface.co/Salesforce/SFR-Embedding-Mistral (Meng et al., 2024), the top 1 model as of May 19, 2024: 1. BAAI/bge-en-icl [^3] https://huggingface.co/BAAI/bge-en-icl (Xiao et al., 2023); 2. dunzhang/stella_en_1.5B_v5444https://huggingface.co/dunzhang/stella_en_1.5B_v5; 3. nvidia/NV-Retriever-v1555https://huggingface.co/nvidia/NV-Retriever-v1 (de Souza et al., 2024); 4. Alibaba-NLP/gte-Qwen2-7B-instruct [^6] https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct (Li et al., 2023); 5. Linq-AI-Research/Linq-Embed-Mistral [^7] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral (Kim et al., 2024); 6. Salesforce/SFR-Embedding-2_R [^8] https://huggingface.co/Salesforce/SFR-Embedding-2_R (Meng* et al., 2024).
MTEB Ranking Grammar Benchmark Ranking Best Performing Prompt Best Performing Instruct Parameters
bge-en-icl 1 6 Term Only Default 7.11B
@@
NV-Retriever-v1 3 7 Wiki Default 7.11B
gte-Qwen2-7B-instruct 4 4 Term Only Specific 7.61B
Linq-Embed-Mistral 5 2 Term Only Specific 7.11B
SFR-Embedding-2_R 6 3 Term Only Specific 7.11B
SFR-Embedding-Mistral 10 1 Wiki Specific 7.11B
Table 2: Rankings on our benchmark, compared to MTEB, and the best performing configurations for the rerankers.
@@
We use two variants of the query: 1. “Dominant word order (Order of Subject, Object, and Verb)” and 2. the Wikipedia summary from the page “Word order” in English.
Apart from state-of-the-art embedding models, we evaluate BM25 itself as the baseline.
Plots with NDCG@k across all values of k for best performing configurations of each retriever on all grammars are presented in Appendix C, and the plots showcasing comparison between two types of prompts are available in Appendix D.
@@
In conclusion, based on the results on the benchmark described in this section, we have selected SFR-Embedding-Mistral with Specific Instruct and Wikipedia Query as the reranker component for the RAG pipeline due to its superior NDCG@20 score compared to other models. Furthermore, the results on the retriever benchmark reinforce the decision to select BM25 as the base retriever, since despite being a rule-based method invented decades ago, it has proven itself to be on a similar level compared to state-of-the-art retrievers on the fragment of the linguistic domain presented in our benchmark.
@@ 5.1 Data
The benchmark for the RAG pipeline comprises 148 descriptive grammars. We selected the ostensibly arbitrary number (initially 150) due to benchmarks with fewer than 100 items being unreliable as a tool of assessment: a wrong answer on one item would result in accuracy decreasing by more than one percentage point. Selecting the grammars in a random fashion would defeat the purpose of the benchmark due to possible biases towards languages having the same descent (a particular family being overrepresented) or being spoken in the same area of the world, without regard to the factual proportions of languages spoken in different areas. Therefore, we used the Genus-Macroarea method described by (Miestamo et al., 2016): in particular, its modification presented by (Cheveleva, 2023). Identically to the method described in (Cheveleva, 2023), we obtain the proportions of languages across macroareas from the list of genera from WALS, automatically choose descriptive grammars from Glottolog’s (Hammarström et al., 2024) References [^9] https://glottolog.org/langdoc database, and create the sample anew, placing the limit of one language for each genus. Our sampling strategy differs from the one utilized by (Cheveleva, 2023) in that we limit our sample to grammars written in English, and instead of restricting references to grammar_sketch and grammar types from Glottolog, we allow references to contain other tags, as long as either grammar_sketch or grammar is present. The resulting proportions of languages across macroareas are present in Table 3.
@@
We annotated four typological features for the languages described by the grammars in the benchmark. The first feature is identical to the one presented in the retriever benchmark: WALS 81A – Order of Subject, Object, and Verb, as an example of a largely self-explanatory and straightforward feature, which is concentrated in one place in the majority of the grammars: if a paragraph mentions which basic constituent order the language has, the mention is in most cases explicit.
@@
The third feature is a complex comprised of 7 binary features all relating to polar (yes/no) questions; we have chosen it in order to evaluate LLMs’ capability to reason on the linguistic domain while taking into account several realizations of the same phenomenon simultaneously. This feature will be further referred to as WALS 116A*: despite being related to WALS 116A, it is more accurately described as an amalgamation of seven separate features from Grambank: GB257, GB260, GB262, GB263, GB264, GB286, and GB291. This feature is essentially a multilabel classification with seven labels: each label/strategy from the set (Interrogative intonation, Interrogative word order, Clause-initial question particle, Clause-medial question particle, Clause-final question particle, Interrogative verb morphology, Tone) is annotated as 1 if it can be used to form polar questions in the described language, and as 0 otherwise.
The last feature is WALS 49A – Number of Cases. It has been chosen due to its quantitative nature, as opposed to binary features determining presence and absence of a particular phenomenon, and its scattered nature: the guideline for WALS 49A takes a liberal approach to determining the number of cases, allowing to consider adpositional clitics as case markers; consequently, the relevant information may be in entirely different sections of the grammar (nominal morphology for “traditional” case markers and syntax for adpositional clitics). Due to Number of Cases being the most time-consuming feature to annotate, we created a separate benchmark for it, consisting partially of grammars from the already existing benchmark and partially of grammars with existing annotations for WALS 49A, maintaining the macroarea proportions and adhering to the one-language-per-genus rule.
@@ 5.2 Results: Evaluating the RAG Pipeline
In order to evaluate any NLP task on a benchmark, it is crucial to run tests in order to determine if the LLM already possesses knowledge about the feature values in the benchmark. It is different from a contamination test: an LLM may possess knowledge about word order in a particular language from Wikipedia or any other Internet resource, while a contamination test would determine if the LLM saw the grammar itself during pretraining.
In order to set the baseline for GPT-4o, i. e. to estimate how well it performs on the linguistic domain without additional information from the grammar, we conducted a test on the RAG pipeline excluding the retrieval module. We prompted GPT-4o to determine the values of all benchmark features based solely on the language name.
feature WALS 81A GB 107 WALS 49A
F1 average micro macro weighted micro macro weighted micro macro weighted
baseline 0,5735 0,2699 0,5534 0,5685 0,5639 0,5700 0,3857 0,1804 0,3145
wikipedia 0,6892 0,7179 0,6694 0,6622 0,5718 0,5957 0,5203 0,4711 0,5314
wikipedia+CoT 0,7027 0,6097 0,6890 0,6554 0,5731 0,5959 0,5405 0,4494 0,5542
reranker 0,7095 0,6141 0,6790 0,6757 0,6011 0,6221 0,5541 0,4332 0,5605
reranker+CoT 0,7297 0,6139 0,6995 0,7095 0,6546 0,6713 0,5135 0,4042 0,5185
feature clause-initial particle clause-medial particle clause-final particle
F1 average micro macro weighted micro macro weighted micro macro weighted
baseline 0,7190 0,5715 0,7398 0,6529 0,4706 0,6426 0,5372 0,5372 0,5379
wikipedia 0,9054 0,7977 0,9054 0,8311 0,7638 0,8439 0,7230 0,7173 0,7314
wikipedia+CoT 0,9189 0,8335 0,9205 0,8851 0,8201 0,$8$ 0,7568 0,7463 0,7644
reranker 0,9459 0,8890 0,9470 0,8378 0,7709 0,8496 0,7027 0,6972 0,7116
reranker+CoT 0,9257 0,8504 0,9278 0,8581 0,7832 0,8641 0,7838 0,7717 0,7902
feature interrog. intonation only interrog. word order interrog. verb morph. tone
F1 average micro macro weighted micro macro weighted micro macro weighted micro macro weighted
baseline 0,4380 0,4038 0,3837 0,9752 0,6937 0,9679 0,7521 0,4292 0,6953 0,9339 0,5828 0,9212
wikipedia 0,8446 0,8392 0,8480 0,9932 0,9427 0,9936 0,7973 0,7075 0,8192 0,9257 0,7407 0,9390
wikipedia+CoT 0,8986 0,8946 0,9007 0,9865 0,8965 0,9878 0,8311 0,7354 0,8451 0,9595 0,8225 0,9637
reranker 0,8378 0,8335 0,8415 0,9865 0,8715 0,9865 0,8243 0,7282 0,8396 0,9662 0,8594 0,9704
reranker+CoT 0,8784 0,8739 0,8810 0,9797 0,8281 0,9808 0,8311 0,7419 0,8465 0,9730 0,8501 0,9730
Table 4: F1 scores across all configurations of the RAG pipeline.
@@
The common components for all prompts (including the baseline prompt without RAG) are listed in Appendix E. All prompts also include the Wikipedia summary for the article about the corresponding feature.

### v2

@@ Abstract
Recent advances in language modeling have demonstrated significant improvements in zero-shot capabilities, including in-context learning, instruction following, and machine translation for extremely under-resourced languages (Tanzer et al., 2024). However, many languages with limited written resources rely primarily on descriptions of grammar and vocabulary.
@@
This set of benchmarks offers the first comprehensive evaluation of language models’ in-context ability to accurately interpret and extract linguistic features, providing a critical resource for scaling NLP to low-resource languages. The code and data are publicly available at https://github.com/al-the-eigenvalue/RAG-on-grammars.
@@ 1 Introduction
The advent of text-based foundational models has accelerated advancements in natural language processing, enhancing multilingual capabilities and applied tasks such as zero-shot machine translation, reading comprehension, and information extraction. Innovations like Machine Translation from One Book (MTOB) (Tanzer et al., 2024) utilize descriptive grammars to improve translation performance for extremely low-resource languages, showing the potential of large-scale language models in bridging linguistic theory and practical NLP applications. Despite the promise, using descriptive grammars for zero-shot MT presents challenges like terminology variability, non-standard structures, and scattered relevant information.
@@
1. The first scaled linguistic evaluation of the LLM machine reading capabilities on descriptive grammars.
2. A pipeline based on Retrieval Augmented Generation (RAG), which extracts relevant paragraphs from grammars based on a given typological characteristic (for example, WALS 81A: Order of Subject, Object and Verb [^1] https://wals.info/feature/81A) and provides them as prompts to an LLM to determine the meaning of these characteristics (for example, Subject-Verb-Object). The pipeline is evaluated through extensive experiments.
3. A benchmark consisting of 700 paragraphs from 14 descriptive grammars, annotated according to whether a linguist can unambiguously determine information about word order (WALS 81A) in the language described, in order to evaluate the quality of information retrieval methods on the task of filtering out the irrelevant paragraphs separately from the RAG pipeline.
4. A benchmark for the RAG pipeline, consisting of 148 grammars for each feature, in order to assess LLMs’ capabilities of determining typological characteristics based on the entire grammar at once and evaluate the effectiveness of different combinations of RAG pipeline components (i.e., information retrieval methods and prompts).
The proposed framework, alongside the presented benchmarks, aims to contribute to the ongoing efforts to improve the quality and efficiency of machine translation systems and to aid linguists in typological research by semi-automating extraction of data from descriptive grammars and. The entire codebase is open-source, licensed under MIT, and available at https://anonymous.4open.science/r/from-MTEB-to-MTOB.
@@ 2 Related Work
Several advances in language modeling outline the possibility to leverage linguistic descriptions: retrieval-augmented generation methods for information extraction and generation, in-context abilities to operate with the descriptive texts in the prompt, and the availability of the texts in the machine-readable format.
@@ 2.1 Retrieval-Augmented Generation
RAG’s potential for handling low-resource languages and complex descriptions, such as linguistic grammars, has not been fully explored. Recent advances in retrieval-based frameworks point toward its applicability in linguistic resource-scarce domains, where available data is often fragmented and incomplete (Gao et al., 2024). This paper leverages the principles of RAG to enhance machine translation and language modeling for under-resourced languages through the use of descriptive grammars.
@@ 2.2 Zero-shot Learning
However, current zero-shot models still struggle with languages that have extremely limited or no monolingual or bilingual corpora. Recent studies, such as (Tanzer et al., 2024) and (Zhang et al., 2024), have shown that integrating linguistic descriptions—such as those found in grammars—can significantly improve zero-shot performance for these languages. In this work, we aim to further this line of research by evaluating the ability of models to utilize descriptive grammars in zero-shot settings.
@@ 4.1 Data
Examples of paragraphs for every relevance category are provided in Appendix B.
Among the resulting 700 paragraphs, 38.86% (annotated with 0) are not relevant to order of any elements in a language. Furthermore, 42.43% of the paragraphs (the ones annotated with 1, 2, 3) are potentially misleading data due to describing order of components in a language without explicitly stating the value of WALS 81A. Finally, 18.71% of paragraphs provide relevant information; however, out of the relevant paragraphs, 31.29% are only partially relevant (annotated with 4), and do not sufficiently elaborate on the constituent order in the monotransitive construction in order for the linguist to be able to determine the value of the feature. More detailed data on the benchmark is provided in Appendix C.
@@ 4.2 Results: Evaluating Rerankers
Wiki Summary 0.7474 0.7601 0.6538 0.7520 0.7690 0.7676 0.7776
Table 1: NDCG@20 scores on the reranker benchmark. Term Only refers to using the sequence “Dominant word order (Order of Subject, Object, and Verb)” as a query, while Wiki Summary refers to using the Wikipedia summary from the page “Word order” in English as a query. The best result for each model is shown in bold, and the best result across all configuration variants is underlined. The NDCG@20 score for BM25 without a reranker is 0.7494.
MTEB Ranking Grammar Benchmark Ranking Best Performing Query Best Performing Instruct Parameters
bge-en-icl (Xiao et al., 2023) 1 6 Term Only Default 7.11B
@@
NV-Retriever-v1 (de Souza et al., 2024) 3 7 Wiki Default 7.11B
gte-Qwen2-7B-instruct (Li et al., 2023) 4 4 Term Only Specific 7.61B
Linq-Embed-Mistral (Kim et al., 2024) 5 2 Term Only Specific 7.11B
SFR-Embedding-2_R (Meng* et al., 2024) 6 3 Term Only Specific 7.11B
SFR-Embedding-Mistral (Meng et al., 2024) 10 1 Wiki Specific 7.11B
Table 2: Rankings on our benchmark, compared to MTEB, and the best performing configurations for the rerankers. The evaluation of the rerankers on our benchmark requires approximately 1.8 GPU hours in total on 1x NVIDIA A100 40GB.
The rerankers we chose are the 6 models with the best NDCG@10 score on the Massive Text Embedding Benchmark (MTEB) leaderboard (Muennighoff et al., 2023) on the Retrieval task for English as of Aug 29, 2024, along with SFR-Embedding-Mistral (Meng et al., 2024), the top 1 model on MTEB as of May 19, 2024. The HuggingFace pages for the models are referenced in Appendix  A.
@@
Apart from the two variants of the instruct, we use two variants of the query: 1. “Dominant word order (Order of Subject, Object, and Verb)” and 2. the Wikipedia summary from the page “Word order” in English.
In addition to state-of-the-art embedding models, we evaluate BM25 itself as the baseline.
Plots with NDCG@k across all values of k for best performing configurations of each reranker on all grammars are presented in Appendix D, and the plots showcasing comparison between two query types are available in Appendix E.
@@
In conclusion, based on the results on the benchmark described in this section, we have selected SFR-Embedding-Mistral with Specific Instruct and Wikipedia Query as the reranker component for the RAG pipeline due to its superior NDCG@20 score compared to other models. Furthermore, the results on the benchmark for rerankers reinforce the decision to select BM25 as the base retriever, since despite being a rule-based method invented decades ago, it has proven itself to be on a similar level compared to state-of-the-art retrievers on the fragment of the linguistic domain presented in our benchmark.
@@ 5.1 Data
The benchmark for the RAG pipeline comprises 148 descriptive grammars. We selected the ostensibly arbitrary number (initially 150) due to benchmarks with fewer than 100 items being unreliable as a tool of assessment: a wrong answer on one item would result in accuracy decreasing by more than one percentage point.
Selecting the grammars in a random fashion would defeat the purpose of the benchmark due to possible biases towards languages having the same descent (a particular family being overrepresented) or being spoken in the same area of the world, without regard to the factual proportions of languages spoken in different areas. Therefore, we used the Genus-Macroarea method described by (Miestamo et al., 2016): in particular, its implementation presented by (Cheveleva, 2023). Identically to the method described in (Cheveleva, 2023), we obtain the proportions of languages across macroareas from the list of genera from WALS, automatically choose descriptive grammars from Glottolog’s (Hammarström et al., 2024) References [^2] https://glottolog.org/langdoc database, and create the sample anew, placing the limit of one language for each genus. Our sampling strategy differs from the one utilized by (Cheveleva, 2023) in that we limit our sample to grammars written in English, and instead of restricting references to grammar_sketch and grammar types from Glottolog, we allow references to contain other tags, as long as either grammar_sketch or grammar is present. The resulting proportions of languages across macroareas are presented in Table 3.
@@
We annotated four typological features for the languages described by the grammars in the benchmark. The first feature is identical to the one presented in the benchmark for rerankers: WALS 81A – Order of Subject, Object, and Verb, as an example of a largely self-explanatory and straightforward feature, which is concentrated in one place in the majority of the grammars: if a paragraph mentions which basic constituent order the language has, the mention is in most cases explicit.
@@
The third feature is a complex composed of seven binary features all related to polar (yes / no) questions; we have chosen it to evaluate the ability of LLMs to reason in the linguistic domain while taking into account several realizations of the same phenomenon simultaneously. This feature will be further referred to as WALS 116A*: despite being related to WALS 116A, it is more accurately described as an amalgamation of seven separate features from Grambank: GB257, GB260, GB262, GB263, GB264, GB286, and GB291. This feature is essentially a multilabel classification with seven labels: each label/strategy from the set (Interrogative intonation, Interrogative word order, Clause-initial question particle, Clause-medial question particle, Clause-final question particle, Interrogative verb morphology, Tone) is annotated as 1 if it can be used to form polar questions in the described language, and as 0 otherwise.
The last feature is WALS 49A - Number of Cases. It has been chosen due to its quantitative nature, as opposed to binary features determining presence and absence of a particular phenomenon, and its scattered nature: the guideline for WALS 49A takes a liberal approach to determining the number of grammatical cases in a language, allowing adpositional clitics to be considered as case markers; consequently, the relevant information may be in entirely different sections of the grammar (nominal morphology for "traditional" case markers and syntax for adpositional clitics). Due to Number of Cases being the most time-consuming feature to annotate, we created a separate benchmark for it, consisting partially of grammars from the already existing benchmark and partially of grammars with existing annotations for WALS 49A, maintaining the macroarea proportions and adhering to the one-language-per-genus rule.
The distribution of values for each feature is presented in Appendix H.
@@ 5.2 Results: Evaluating the RAG Pipeline
In order to evaluate any NLP task on a benchmark, it is crucial to run tests in order to determine if the LLM already possesses knowledge about the feature values in the benchmark. It is different from a contamination test: an LLM may possess knowledge about word order in a particular language from any resource, while a contamination test would determine if the LLM saw the grammar itself during pretraining.
In order to set the baseline for GPT-4o, i.e., to estimate how well it performs on the linguistic domain without additional information from the grammar, we conducted a test on the RAG pipeline excluding the retrieval module. We prompted GPT-4o to determine the values of all benchmark features without the retrieved paragraphs: the only information given to the model consisted of the prompt and the Wikipedia summary for the article about the corresponding feature. The prompts are listed in Appendix F. For each feature, the baseline run was executed ten times in order to capture the variance of the results and to represent presence vs. absence of knowledge more accurately. It is crucial to note that the baseline was evaluated only on those languages where the grammar itself had sufficient information to identify the feature: 136, 146, 121, and 140 values out of 148 for WALS 81A, GB 107, WALS 116A, and WALS 49A respectively.
F1 mean Baseline BM25 BM25+CoT Reranker Rer.+CoT
micro 0.5551 ± 0.0359 0.6892 0.7027 0.7095 0.7297
macro 0.2812 ± 0.0276 0.7179 0.6097 0.6141 0.6139
WALS 81A weighted 0.5328 ± 0.0337 0.6694 0.6890 0.6790 0.6995
micro 0.5747 ± 0.0350 0.6622 0.6554 0.6757 0.7095
macro 0.5740 ± 0.0354 0.5718 0.5731 0.6011 0.6546
GB 107 weighted 0.5724 ± 0.0361 0.5957 0.5959 0.6221 0.6713
micro 0.3986 ± 0.0238 0.5203 0.5405 0.5541 0.5135
macro 0.2216 ± 0.0158 0.4711 0.4494 0.4332 0.4042
WALS 49A weighted 0.3453 ± 0.0237 0.5314 0.5542 0.5605 0.5185
micro 0.4504 ± 0.0396 0.8446 0.8986 0.8378 0.8784
macro 0.4240 ± 0.0456 0.8392 0.8946 0.8335 0.8739
interrog. intonation only weighted 0.4068 ± 0.0484 0.8480 0.9007 0.8415 0.8810
micro 0.9653 ± 0.0076 0.9932 0.9865 0.9865 0.9797
macro 0.6641 ± 0.0460 0.9427 0.8965 0.8715 0.8281
interrog. word order weighted 0.9611 ± 0.0064 0.9936 0.9878 0.9865 0.9808
micro 0.7264 ± 0.0402 0.9054 0.9189 0.9459 0.9257
macro 0.5446 ± 0.0502 0.7977 0.8335 0.8890 0.8504
clause- initial particle weighted 0.7371 ± 0.0337 0.9054 0.9205 0.9470 0.9278
micro 0.4909 ± 0.0409 0.7230 0.7568 0.7027 0.7838
macro 0.4808 ± 0.0415 0.7173 0.7463 0.6972 0.7717
clause- final particle weighted 0.4661 ± 0.0428 0.7314 0.7644 0.7116 0.7902
micro 0.6760 ± 0.0367 0.8311 0.8851 0.8378 0.8581
macro 0.5011 ± 0.0544 0.7638 0.8201 0.7709 0.7832
clause- medial particle weighted 0.6644 ± 0.0355 0.8439 0.8888 0.8496 0.8641
micro 0.7678 ± 0.0185 0.7973 0.8311 0.8243 0.8311
macro 0.4533 ± 0.0311 0.7075 0.7354 0.7282 0.7419
interrog. verb morphology weighted 0.7102 ± 0.0160 0.8192 0.8451 0.8396 0.8465
micro 0.9273 ± 0.0128 0.9257 0.9595 0.9662 0.9730
macro 0.5147 ± 0.0573 0.7407 0.8225 0.8594 0.8501
tone weighted 0.9104 ± 0.0121 0.9390 0.9637 0.9704 0.9730
Table 4: F1 scores across all configurations of the pipeline. All five configurations contain prompts from Appendix F with Wikipedia summaries for corresponding features. The Baseline column refers to prompting GPT-4o without materials from grammars (i.e, without RAG). BM25 refers to using 50 paragraphs retrieved by BM25. Reranker refers to using 20 paragraphs selected from these 50 by SFR-Embedding-Mistral. CoT refers to Chain-of-Thought (adding instructions and examples from WALS or Grambank to the prompt). The best F1 value in each row is shown in bold. The values after the ± sign in the Baseline column are the sample standard deviation across 10 runs. Each of the remaining four configurations was executed only once.
@@
More details regarding the prompt configurations are available in Appendix G.
The temperature value we chose for GPT-4o is 0.2, since our tasks requires the model to incline towards more deterministic behavior.
Running the rerankers on the RAG benchmark for all configurations requires approximately 7.2 GPU hours in total on 1x NVIDIA A100 40GB.